expand
(X^3-1/X^2)^6

Answer by Boreal(15235)

I will write as x^3-x^(-2) all raised to the 6th power
The coefficients will be 1,6,15,20,15,6,1
The first term is x^18
The second term is 6*x^15*x^(-2)=6x^13
The third term is 15*(x^12)*x^(-4)=15x^8
The fourth term is 20(x^9)*x^(-6)=20x^3
The fifth term is 15 (x^6)*(x^(-8)=15x^(-2)
The sixth term is 6 (x^3)*(x^(-10)=6/x^7
the last term is x^(-12)=1/x^12
the expansion is the sum of all of these.